OpenStudy (anonymous):

Find the derivative of (x^2+1)^x.

zepdrix (zepdrix):

Ooo this one is a stinker!! We have our variable term in the base AND in the exponent. So neither the power rule, nor the rule for exponents will work. We should use logs to get through this.

OpenStudy (anonymous):

take log and then diff.

OpenStudy (anonymous):

Can you show the work?

zepdrix (zepdrix):

\[\Large\bf\sf y=(x^2+1)^x\qquad\to\qquad \ln y=\ln(x^2+1)^x\]Remember this rule for logs?\[\Large\bf\sf \color{orangered}{\log(a^b)\quad=\quad b\cdot \log(a)}\]

OpenStudy (anonymous):

So ln y=x*ln(x^2+1)?

zepdrix (zepdrix):

good good.

zepdrix (zepdrix):

Now differentiate (with respect to x), applying the product rule on the right side.

OpenStudy (anonymous):

But on the left side, it will be (1/y)(dy/dx), right?

zepdrix (zepdrix):

Yes, good :)

OpenStudy (anonymous):

Okay, so I got (1/y)(dy/dx)=2x^2/(x^2+1)+ln(x^2+1). But how do I simplify this by multiply y to the right side? What will it look like?

zepdrix (zepdrix):

Ya multiply y to the other side, then we want to substitute our original equation into that y.\[\Large\bf\sf y=(x^2+1)^x\]

OpenStudy (anonymous):

So how does it look like?

zepdrix (zepdrix):

We want our derivative in terms of `x` alone.

zepdrix (zepdrix):

\[\Large\bf\sf y'=(x^2+1)^x\left[\ln(x^2+1)+\frac{2x^2}{x^2+1}\right]\]

zepdrix (zepdrix):

That is a good place to stop. Good looking final answer. You COULD find a common denominator and then it simplifies down a little bit further. But it's not really necessary.

OpenStudy (anonymous):

I got it. Exactly the answer I was looking for. Thanks a lot.

zepdrix (zepdrix):

Cool \c:/ good job